Given the following data, perform the appropriate calculations and evaluate the values related to oxygen content and oxygen consumption. PaO2 = 65 mm Hg Hb = 10 g/dl SaO2 = 91% PVO2 = 30 mm Hg SVO2 = 50% cardiac output = 16 l/min

CaO2 = 12.39cc O2/dl

      -       A.       B.   

a-v DO2 = 5.6 cc O2/dl

      -       A.       B.   

This patient is at rest

      -       A.       B.   

VO2 = 89.6 cc/min
Answer #1

ANSWER

    * a -v DO2 = 5.6 cc O2/dl. ( A . Correct )

            Formula = CaO2 - CvO2

               CaO2 = 12.39 cc O2/dl.

               CvO2 = (Hb * 1.34 * SVO2) + (PaO2 * 0.003)

                     Hb = 10 gm /dl.

                 SVO2 = 50%.(0.50gm)

               PaO2= 65 mmHg.

                          = (10 * 1.34 * O.50) + (65 * 0.003).

                          = 6.7 + 0.195

                          = 6.89 cc O2/dl.

          a - v DO2 = 12.39 - 6.89

                          = 5.5 cc O2/dl.

* This patient is at rest. ( A. Correct ).

                * Arteriovenous oxygen content difference is calculated through the the amount of oxygen

                  taken from blood by the tissues.in resting period it 5ml in every 100ml is extracted and

                  produce a - v DO2 of 5 cc O2/dl.

* VO2 = 89.6 cc / min.( B. Incorrect)

          Formula = Cardiac output * (CaO2 - CvO2)

               Cardiac output = 16 l /min.

                  CaO2 - CvO2 = 5.5 cc O2 /dl.

                                          = 16 * 5.5

                                         = 88 cc /min